# Fullness of $q$-Araki-Woods Factors

### v1

@@ Abstract.
The $q$-Araki-Woods factor associated to a group of orthogonal transformations on a real Hilbert space $\mathsf{H}_{\mathbb{R}}$ is full as soon as $\dim\mathsf{H}_{\mathbb{R}}\geq 2$.
@@ 1. Introduction
[…]
Recall that the $q$-Araki-Woods factors, denoted as $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$, are associated with a strongly continuous one-parameter group of orthogonal transformations $\{U\}_{t\in\mathbb{R}}$ on a real Hilbert space $\mathsf{H}_{\mathbb{R}}$ (see the construction in Section 2). These von Neumann algebras were introduced by Hiai [Hia], combining the quasi-free deformations of Shlyakhtenko [Shl] and $q$-deformations of Bożejko-Speicher [BS]. While various structural properties are known for the free Araki-Woods factors $\Gamma_{0}(\mathsf{H}_{\mathbb{R}},U_{t})$, such as fullness, solidity, strong solidity, the absence of Cartan subalgebras, and complete metric approximation properties [Shl, HRau, HRic, BHV], many of these properties remain elusive for non-zero $q$, with only a few of these results available [ABW, BM, BMRW, KSW, Nou, SW]. For instance, the fundamental question of factoriality of $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$ was only recently addressed in complete generality in [KSW], with some major success achieved in [BMRW]. Furthermore, the factors $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$ are known to be non-injective [KSW, Nou]. Drawing inspiration from the free/undeformed case, it is then natural to ask whether $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$ is full as well for non-zero $q$. In this paper, we consider this problem and prove the following result.
@@
We provide a concise overview of previous attempts and highlight our contribution to the proof. So far, there are mainly two papers, [KSW] and [HI], which have dealt with the fullness of $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$ for $q\neq 0$. In order to describe them, it is crucial to first mention that the Hilbert space $\mathsf{H}_{\mathbb{R}}$ can be decomposed into direct sums of almost periodic $\mathsf{H}_{\mathbb{R}}^{ap}$ and weakly mixing $\mathsf{H}_{\mathbb{R}}^{wm}$ subspaces, where $\mathsf{H}_{\mathbb{R}}^{ap}$ is the direct sum of all finite-dimensional sub-representations of $U_{t}$, and $\mathsf{H}_{\mathbb{R}}^{wm}$ encompasses the rest. When $\mathsf{H}_{\mathbb{R}}$ is finite-dimensional, $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$ is shown to be full [KSW]. The proof involves establishing the non-injectivity of $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$ by combining results on conjugate variables and abstract findings of Nelson [Nel]. It then proceeds to demonstrate Ozawa’s AO property for $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$ before invoking the abstract results of Houdayer-Raum [HRau] to conclude fullness. In contrast, Houdayer and Isono [HI] show that if $\mathsf{H}_{\mathbb{R}}^{wm}\neq 0$ or if the point spectrum consisting of eigenvalues of $A$ is infinite and bounded, then $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$ is full. Here, $A$ denotes the analytic generator of the one-parameter group of unitaries $U_{t}$, extended linearly to the complexification of $\mathsf{H}_{\mathbb{R}}$. Their proof follows the ideas of Śniady [Sni] from the tracial case, i.e. when $U_{t}$ is trivial. However, their approach encounters difficulties when there is a sequence of eigenvalues growing sufficiently fast to infinity. In this paper, we overcome such challenges by adopting a distinct approach.
@@
Our arguments involve a meticulous evaluation of certain Wick products in the centralizer of $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$ associated with vectors in a $2\times 2$ block of $\mathsf{H}_{\mathbb{R}}$. As in [BMRW], we make a crucial use of the non-triviality of the group $(U_{t})$; however, it is worth noting that the methods in [BMRW] only work for the weak convergence of the concerning sequences, which does not apply to the study of fullness. In our approach, we establish several ’norm’-convergence results involving the Wick products and creation/annihilation operators. The key novelty in these convergence results lies in the utilization of an ‘averaging technique’, where one will observe that the sequences themselves may fail to converge. For instance, when $\xi$ and $\eta$ are two unit vectors in $\mathsf{H}_{\mathbb{R}}$, the bounded sequence $T_{n}=(1-q)^{n}c(\xi)^{n}c(\eta)^{n}$ never converges in norm, although it does converge to 00 in weak operator topology, where $c(\xi)$ denotes the creation operator on the (deformed) Fock space. Nonetheless, we demonstrate that the Cesaro sum of $T_{n}$ converges to 00 in norm with the additional assumption that $\xi,\eta$ are orthogonal to each other (see Proposition 3.4). The main idea then is to carefully select such vectors $\xi$ and $\eta$, and employ these convergence results to prove fullness (see Theorem 1.1). It is crucial to emphasize that the methods employed in this paper are not applicable to the cases considered in the previous two attempts [KSW, HI]. Thus, a combination of both [KSW] and [HI] is necessary to conclusively establish Theorem 1.1 in full generality.
@@ 2. Preliminaries on $q$-Araki-Woods factors
We briefly describe construction of the $q$-Araki-Woods factors and refer the readers to [Hia] for a detailed treatment. As a convention, all Hilbert spaces are separable and inner products are linear in the second variable. For any real or complex Hilbert space $H$, $\mathbf{B}(H)$ denotes the algebra of bounded linear operators on $H$.
Let $\mathsf{H}_{\mathbb{R}}$ be a real Hilbert space, and let $U_{t}$ be a strongly continuous orthogonal representations of $\mathbb{R}$ on $\mathsf{H}_{\mathbb{R}}$. Consider the complexification $\mathsf{H}_{\mathbb{C}}=\mathsf{H}_{\mathbb{R}}\otimes_{\mathbb{R}}\mathbb{C}$ of $\mathsf{H}_{\mathbb{R}}$, whose inner product is denoted by $\langle\cdot,\cdot\rangle_{\mathsf{H}_{\mathbb{C}}}$. Identify $\mathsf{H}_{\mathbb{R}}$ inside $\mathsf{H}_{\mathbb{C}}$ by $\mathsf{H}_{\mathbb{R}}\otimes 1$, so that $\mathsf{H}_{\mathbb{C}}=\mathsf{H}_{\mathbb{R}}+i\mathsf{H}_{\mathbb{R}}$. Given $\xi=\xi_{1}+i\xi_{2}\in\mathsf{H}_{\mathbb{C}}$ for some $\xi_{1},\xi_{2}\in\mathsf{H}_{\mathbb{R}}$, write the canonical involution $\bar{\xi}$ as
@@
We extend $U_{t}$ to a strongly continuous one parameter group of unitaries on the complexification $\mathsf{H}_{\mathbb{C}}$, which is still denoted by $U_{t}$. Let $A$ denote the positive, self-adjoint, and non-singular analytic (unbounded) operator on $\mathsf{H}_{\mathbb{C}}$ obtained via Stone’s theorem, which satisfies $U_{t}=A^{it}$ for all $t\in\mathbb{R}$. Note that $\overline{A\xi}=A^{-1}\bar{\xi}$ for any $\xi$ in the domain of $A$. In particular, $\sigma(A)\cap(0,\infty)$ is symmetric around $1$ i.e. for $\lambda\neq 0$, $\lambda\in\sigma(A)$ if and only if $1/\lambda\in\sigma(A)$, where $\sigma(A)$ is the spectrum of $A$.
@@
$\langle\xi,\zeta\rangle_{U}=\langle 2(1+A^{-1})^{-1}\xi,\zeta\rangle\quad\forall\xi,\zeta\in H_{\mathbb{C}}.$
@@
for all $\xi_{i},\zeta_{i}\in\mathsf{H}_{U}$, $1\leq i\leq n$. Here $S_{n}$ denotes the symmetric group on $n$ elements and $|\pi|$ denotes the number of inversions for the permutation $\pi\in S_{n}$. That the above assignment yields a genuine inner product was proved in [BS]. On $\mathcal{F}_{q}(\mathsf{H}_{U})$, we define for each $\xi\in\mathsf{H}_{U}$ a bounded operator $c(\xi)$ (called left q-creation) and its adjoint $c^{*}(\xi)$ (called left q-annihilation) by:
@@
for all $\xi_{1}\otimes\ldots\otimes\xi_{n}\in\mathsf{H}_{U}^{\otimes_{q}^{n}}$. Here $\hat{\xi_{i}}$ notates the omitted $i$-th letter from the tensor. That $c(\xi)$ and $c^{*}(\xi)$ are bounded operators with them being adjoint to each other is a well known fact (see [BS]). In a similar fashion, one defines the right $q$-creation and right $q$-annihilation operators $r(\xi)$ and $r^{*}(\xi)$ respectively for all $\xi\in\mathsf{H}_{U}$ as follows:
@@
for all $\xi_{1}\otimes\ldots\otimes\xi_{n}\in\mathsf{H}_{U}^{\otimes_{q}^{n}}$. Again $r(\xi)$ and $r^{*}(\xi)$ are bounded operators such that they are adjoint to each other. The creation and annihilation operators satisfy the following $q$-commutation relations:
@@
(2) $r^{*}(\xi)r(\eta)-qr(\eta)r^{*}(\xi)=\langle\xi,\eta\rangle_{U}$
for any $\xi,\eta\in\mathsf{H}_{U}$. We shall repeatably use the above commutation relations without referring to them. In particular, the most useful use of these relations will be the following consequence:
@@
We now consider the von Neumann algebra introduced by Hiai [Hia] (and by Shlyakhtenko [Shl]), denoted $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$, to be the von Neumann subalgebra of ${B}(\mathcal{F}_{q}(\mathsf{H}_{U}))$ generated by the set $\{W(\xi);\xi\in\mathsf{H}_{\mathbb{R}}\}$, where
@@
The algebra $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$ is called the $q$-deformed Araki-Woods von Neumann algebra. If the representation $U_{t}$ is clear from the context, we shall denote the algebra $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$ simply by $\mathsf{M}_{q}$. Since it is now known that $\mathsf{M}_{q}$ is factor i.e. $\mathsf{M}_{q}$ has trivial center (see [KSW]), we will call them $q$-Araki-Woods factors.
The vacuum vector $\Omega$ is cyclic and separating for the factor $\mathsf{M}_{q}$, and hence induces a faithful normal state: $\varphi(\cdot)=\langle\Omega,\cdot\;\Omega\rangle_{q}$ on $\mathsf{M}_{q}$ and is called the $q$-quasi free state. It is immediate to verify that the GNS Hilbert space associated to $\varphi$ is $\mathcal{F}_{q}(\mathsf{H}_{U})$. Since $\Omega$ is a separating vector for both $\mathsf{M}_{q}$ and $\mathsf{M}_{q}^{\prime}$, there exist for each $\xi\in\mathsf{M}_{q}\Omega$ and $\zeta\in\mathsf{M}_{q}^{\prime}\Omega$ unique operators $x_{\xi}\in\mathsf{M}_{q}$ and $x^{\prime}_{\zeta}\in\mathsf{M}_{q}^{\prime}$ such that $\xi=x_{\xi}\Omega$ and $\zeta=x_{\zeta}^{\prime}\Omega$. Write
@@
Note that the notation $W(\xi)$ is in line with the one introduced above for $\xi\in\mathsf{H}_{\mathbb{R}}$. The operators $W(\xi)$ and $W(\zeta)$ are respectively called left and right Wick operators. It is known that $\xi_{1}\otimes\ldots\otimes\xi_{n}\in\mathsf{M}_{q}\Omega$ for all $\xi_{i}\in\mathsf{H}_{\mathbb{C}},1\leq i\leq n$ and $\zeta_{1}\otimes\ldots\otimes\zeta_{m}$ for all $\zeta_{j}\in\mathsf{H}_{\mathbb{C}}\cap\operatorname{Dom}(A^{-1/2}),1\leq j\leq m$. By definition, the operators $\mathsf{W}(\xi)$ and $W_{r}(\zeta)$ commute for all $\xi\in\mathsf{M}_{q}\Omega$ and $\zeta\in\mathsf{M}_{q}^{\prime}\Omega$. In fact, the commutant of $\mathsf{M}_{q}$ in $B(\mathcal{F}_{q}(\mathsf{H}_{U}))$ is given by
@@
For any vector $\xi_{1},\ldots,\xi_{n}\in\mathsf{H}_{\mathbb{R}}+i\mathsf{H}_{\mathbb{R}}$, the Wick-formula for operators in $\mathsf{M}_{q}$ is given by
@@
For any vector $\xi_{1},\ldots,\xi_{n}\in\mathsf{H}_{\mathbb{R}}^{\prime}+i\mathsf{H}_{\mathbb{R}}^{\prime}$, the Wick-formula for operators in $\mathsf{M}_{q}^{\prime}$ is given by
@@
The following lemmas are either well-known or easy to verify.
@@
$c^{*}(f)(\xi\otimes\eta)=(c^{*}(f)\xi)\otimes\eta+q^{n}\xi\otimes c^{*}(f)\eta$
@@
We end this section by recalling the general structure of a one parameter group of orthogonal operators. Let $\{U_{t}\}_{t\in\mathbb{R}}$ be a strongly continuous orthogonal representation of $\mathbb{R}$ on a real Hilbert space $\mathsf{H}_{\mathbb{R}}$ and let $A$ denote the analytic generator of the complexification of $U_{t}$. Let $\mathsf{H}_{\mathbb{R}}^{ap}$ denote the closed real subspace generated by the eigenvectors of $A$ and let $\mathsf{H}_{\mathbb{R}}^{wm}={\mathsf{H}_{\mathbb{R}}^{ap}}^{\perp}$. The subspaces $\mathsf{H}_{\mathbb{R}}^{ap}$ and $\mathsf{H}_{\mathbb{R}}^{wm}$ are reducing under $A$ and every $U_{t}$, and they are respectively called the almost periodic and weakly mixing part of $\mathsf{H}_{\mathbb{R}}$. Similarly, ${U_{t}}_{|_{\mathsf{H}_{\mathbb{R}}^{ap}}}$ and ${U_{t}}_{|_{\mathsf{H}_{\mathbb{R}}^{wm}}}$ are called the almost periodic and weakly mixing part of $U_{t}$ respectively.

### v2

@@ Abstract.
The $q$-Araki-Woods factor associated to a group of orthogonal transformations on a real separable Hilbert space $\mathsf{H}_{\mathbb{R}}$ is full as soon as $\dim\mathsf{H}_{\mathbb{R}}\geq 2$.
@@ 1. Introduction
A von Neumann factor $\mathsf{M}$ is called full if, for any bounded net ${x_{i}}$ in $\mathsf{M}$ such that $\lim_{i}\|\varphi(\cdot x_{i})-\varphi(x_{i}\cdot)\|_{\mathsf{M}_{*}}=0$ for all $\varphi\in\mathsf{M}_{*}$, there exists a bounded net ${\lambda_{i}}$ of scalars such that $\lim_{i}(x_{i}-\lambda_{i})=0$-strongly [Con1]. Here, $\mathsf{M}_{*}$ denotes the predual of $\mathsf{M}$ i.e. the space of normal linear functionals on $\mathsf{M}$. A full factor is always non-injective or non-hyperfinite i.e. it cannot be approximated by finite-dimensional subalgebras. For separable type II1 factors, the negation of fullness is equivalent to Murray and von Neumann’s notion of Property Gamma [Con2]. Connes established several other equivalent conditions for fullness in type II1 factors in his seminal work [Con2] on the classification program of injective factors. Some of these equivalent conditions, such as the spectral gap property, do not seem to generalize easily to type III factors. However, in a major breakthrough recently, Marrakchi [Mar] proved that fullness for separable type III1 factors is equivalent to the assertion that the $C^{*}$-algebra generated by $\mathsf{M}$ and its commutant $\mathsf{M}^{\prime}$, represented in the standard form, contains compact operators. The focus of this paper is to utilize one of these equivalent conditions to establish, in full generality, that the $q$-Araki-Woods factors are full.
Recall that the $q$-Araki-Woods factors, denoted as $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$, are associated with a number $q\in(-1,1)$ and a strongly continuous one-parameter group of orthogonal transformations $\{U_{t}\}_{t\in\mathbb{R}}$ on a real Hilbert space $\mathsf{H}_{\mathbb{R}}$ (see the construction in Section 2). These von Neumann algebras were introduced by Hiai [Hia], combining the quasi-free deformations of Shlyakhtenko [Shl] and $q$-deformations of Bożejko-Speicher [BS]. While various structural properties are known for the free Araki-Woods factors $\Gamma_{0}(\mathsf{H}_{\mathbb{R}},U_{t})$, such as fullness, solidity, strong solidity, the absence of Cartan subalgebras, and complete metric approximation properties [Shl, HRau, HRic, BHV], many of these properties remain elusive for non-zero $q$, with only a few of the relevant results available [ABW, BM, BMRW, KSW, Nou, SW]. For instance, the fundamental question of factoriality of $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$ was only recently addressed in complete generality in [KSW], with some major success achieved earlier in [BMRW]. Furthermore, the factors $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$ are known to be non-injective [KSW, Nou]. Drawing inspiration from the free/undeformed case, it is then natural to ask whether $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$ is full as well for non-zero $q$. This problem has been considered earlier in [Sni, HI, KSW] with results obtained in various specific cases. In this paper, we address the remaining unresolved cases and hence present the following result in full generality.
@@
We provide a concise overview of previous attempts and highlight our contribution to the proof. When the representation $\{U_{t}\}_{t\in\mathbb{R}}$ is trivial and $q\neq 0$, Śniady [Sni] proved the fullness of $\Gamma_{q}(\mathsf{H}_{\mathbb{R}})$ for a large-dimensional real Hilbert space $\mathsf{H}_{\mathbb{R}}$, dependent on the value of $q$ (including the case when $\dim\mathsf{H}_{\mathbb{R}}=\infty$); additionally, Miyagawa-Speicher [MS] established the fullness of $\Gamma_{q}(\mathsf{H}_{\mathbb{R}})$ for a finite-dimensional Hilbert space $\mathsf{H}_{\mathbb{R}}$ with $\dim\mathsf{H}_{\mathbb{R}}\geq 2$. On the other hand, there are mainly two papers, [KSW] and [HI], which have dealt with the fullness of $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$ for $q\neq 0$ and $\{U_{t}\}_{t\in\mathbb{R}}$ non-trivial. In order to describe them, it is crucial to first mention that the Hilbert space $\mathsf{H}_{\mathbb{R}}$ can be decomposed into direct sums of almost periodic $\mathsf{H}_{\mathbb{R}}^{ap}$ and weakly mixing $\mathsf{H}_{\mathbb{R}}^{wm}$ subspaces, where $\mathsf{H}_{\mathbb{R}}^{ap}$ is the direct sum of all finite-dimensional invariant subspaces of $\{U_{t}\}_{t\in\mathbb{R}}$, and $\mathsf{H}_{\mathbb{R}}^{wm}$ encompasses the rest. When $\mathsf{H}_{\mathbb{R}}$ is finite-dimensional with $\dim\mathsf{H}_{\mathbb{R}}\geq 2$, $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$ is shown to be full in [KSW]. The proof involves establishing the non-injectivity of $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$ by combining results on conjugate variables and general findings of Nelson [Nel]. It then proceeds to demonstrate Ozawa’s AO property for $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$ before invoking the general results of Houdayer-Raum [HRau] to conclude fullness. In contrast, Houdayer and Isono [HI] show that if $\mathsf{H}_{\mathbb{R}}^{wm}\neq 0$ or if the point spectrum, the set of of eigenvalues of $A$, is infinite and bounded, then $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$ is full. Here, $A$ denotes the analytic generator of the one-parameter group of unitaries $\{U_{t}\}_{t\in\mathbb{R}}$, extended linearly to the complexification of $\mathsf{H}_{\mathbb{R}}$. Their proof follows the ideas of [Sni] from the tracial case, i.e. when $\{U_{t}\}_{t\in\mathbb{R}}$ is trivial. However, their approach encounters difficulties when there is a sequence of eigenvalues growing sufficiently fast to infinity. In this paper, we overcome such challenges by adopting a different approach.
@@
Our arguments involve a meticulous evaluation of certain Wick products in the centralizer of $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$ associated with vectors in a $2\times 2$ block of $\mathsf{H}_{\mathbb{R}}$. As in [BMRW], we make a crucial use of the non-triviality of the group $\{U_{t}\}_{t\in\mathbb{R}}$; however, it is worth noting that the methods in [BMRW] only work for the weak convergence of the concerning sequences, which does not apply to the study of fullness. In our approach, we establish several norm convergence results involving the Wick products and creation/annihilation operators. The key novelty in these convergence results lies in the utilization of an ‘averaging technique’, where one will observe that the sequences themselves may fail to converge. For instance, when $\xi$ and $\eta$ are two unit vectors in $\mathsf{H}_{\mathbb{R}}$, the bounded sequence $T_{n}=(1-q)^{n}c(\xi)^{n}c(\eta)^{n}$ never converges in norm, although it does converge to 00 in weak operator topology, where $c(\xi)$ denotes the creation operator on the (deformed) Fock space. Nonetheless, we demonstrate that the Cesáro sum of $T_{n}$ converges to 00 in norm with the additional assumption that $\xi,\eta$ are orthogonal to each other (see Proposition 3.4). The main idea then is to carefully select such vectors $\xi$ and $\eta$, and employ these convergence results to prove fullness (see Theorem 1.1). It is crucial to emphasize that the methods employed in this paper are not applicable to the cases considered in the previous two attempts [KSW, HI]. Thus, a combination of both [KSW] and [HI] is necessary to conclusively establish Theorem 1.1 in full generality.
@@ 2. Preliminaries on $q$-Araki-Woods factors
We briefly describe construction of the $q$-Araki-Woods factors and refer the readers to [Hia] for a detailed treatment. As a convention, all Hilbert spaces are separable and inner products are linear in the second variable. For any real or complex Hilbert space $H$, ${B}(H)$ denotes the algebra of bounded linear operators on $H$.
Let $\mathsf{H}_{\mathbb{R}}$ be a real Hilbert space, and let $\{U_{t}\}_{t\in\mathbb{R}}$ be a strongly continuous orthogonal representation of $\mathbb{R}$ on $\mathsf{H}_{\mathbb{R}}$. Consider the complexification $\mathsf{H}_{\mathbb{C}}=\mathsf{H}_{\mathbb{R}}\otimes_{\mathbb{R}}\mathbb{C}$ of $\mathsf{H}_{\mathbb{R}}$, whose inner product is denoted by $\langle\cdot,\cdot\rangle_{\mathsf{H}_{\mathbb{C}}}$. Identify $\mathsf{H}_{\mathbb{R}}$ inside $\mathsf{H}_{\mathbb{C}}$ by $\mathsf{H}_{\mathbb{R}}\otimes 1$, so that $\mathsf{H}_{\mathbb{C}}=\mathsf{H}_{\mathbb{R}}+i\mathsf{H}_{\mathbb{R}}$. Given $\xi=\xi_{1}+i\xi_{2}\in\mathsf{H}_{\mathbb{C}}$ for some $\xi_{1},\xi_{2}\in\mathsf{H}_{\mathbb{R}}$, write the canonical involution $\bar{\xi}$ as
@@
We extend $\{U_{t}\}_{t\in\mathbb{R}}$ to a strongly continuous one parameter group of unitaries on the complexification $\mathsf{H}_{\mathbb{C}}$, which is still denoted by $\{U_{t}\}_{t\in\mathbb{R}}$. Let $A$ denote the positive, self-adjoint, and non-singular (unbounded) operator on $\mathsf{H}_{\mathbb{C}}$ obtained via Stone’s theorem, which satisfies $U_{t}=A^{it}$ for all $t\in\mathbb{R}$. Note that $\overline{A\xi}=A^{-1}\bar{\xi}$ for any $\xi$ in the domain of $A$. In particular, $\sigma(A)\cap(0,\infty)$ is symmetric around $1$ i.e. for $\lambda\neq 0$, $\lambda\in\sigma(A)$ if and only if $1/\lambda\in\sigma(A)$, where $\sigma(A)$ is the spectrum of $A$.
@@
$\langle\xi,\zeta\rangle_{U}=\langle 2(1+A^{-1})^{-1}\xi,\zeta\rangle\quad\forall\xi,\zeta\in\mathsf{H}_{\mathbb{C}}.$
@@
for all $\xi_{i},\zeta_{i}\in\mathsf{H}_{U}$, $1\leq i\leq n$. Here $S_{n}$ denotes the symmetric group on $n$ elements and $|\pi|$ denotes the number of inversions for the permutation $\pi\in S_{n}$. That the above assignment yields a genuine inner product was proved in [BS]. On $\mathcal{F}_{q}(\mathsf{H}_{U})$, we define for each $\xi\in\mathsf{H}_{U}$ operators $c(\xi)$ (called left q-creation) and $c^{*}(\xi)$ (called left q-annihilation) by the following assignments:
@@
for all $\xi_{1}\otimes\ldots\otimes\xi_{n}\in\mathsf{H}_{U}^{\otimes_{q}^{n}}$. Here $\hat{\xi_{i}}$ notates the omitted $i$-th letter from the tensor. That $c(\xi)$ and $c^{*}(\xi)$ extend to bounded linear operators on $\mathcal{F}_{q}(\mathsf{H}_{U})$ being adjoint to each other is a well known fact (see [BS]). In a similar fashion, one defines the right $q$-creation and right $q$-annihilation operators $r(\xi)$ and $r^{*}(\xi)$ respectively for all $\xi\in\mathsf{H}_{U}$ as follows:
@@
for all $\xi_{1}\otimes\ldots\otimes\xi_{n}\in\mathsf{H}_{U}^{\otimes_{q}^{n}}$. Again $r(\xi)$ and $r^{*}(\xi)$ are bounded operators which are adjoint to each other. The creation and annihilation operators satisfy the following $q$-commutation relations:
@@
$r^{*}(\xi)r(\eta)-qr(\eta)r^{*}(\xi)=\langle\xi,\eta\rangle_{U}$
for any $\xi,\eta\in\mathsf{H}_{U}$. We shall repeatedly use the above commutation relations without referring to them. In particular, the most useful use of these relations will be the following consequence:
@@
We now consider the von Neumann algebra introduced by Hiai [Hia] (following Shlyakhtenko [Shl] and Bożejko-Speicher [BS]), denoted $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$, to be the von Neumann subalgebra of ${B}(\mathcal{F}_{q}(\mathsf{H}_{U}))$ generated by the set $\{W(\xi);\xi\in\mathsf{H}_{\mathbb{R}}\}$, where
@@
The algebra $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$ is called the $q$-deformed Araki-Woods von Neumann algebra. If the representation $\{U_{t}\}_{t\in\mathbb{R}}$ is clear from the context, we shall denote the algebra $\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$ simply by $\mathsf{M}_{q}$. We shall always assume that $\dim\mathsf{H}_{\mathbb{R}}\geq 2$. Since it is now known that $\mathsf{M}_{q}$ is a factor i.e. $\mathsf{M}_{q}$ has trivial center (see [KSW, Hia]), we will call it a $q$-Araki-Woods factor.
The vacuum vector $\Omega$ is cyclic and separating for the factor $\mathsf{M}_{q}$, and hence induces a faithful normal state: $\varphi(\cdot)=\langle\Omega,\cdot\;\Omega\rangle_{q}$ on $\mathsf{M}_{q}$, called the $q$-quasi free state. It is immediate to verify that the GNS Hilbert space associated to $\varphi$ is $\mathcal{F}_{q}(\mathsf{H}_{U})$. Since $\Omega$ is a separating vector for both $\mathsf{M}_{q}$ and $\mathsf{M}_{q}^{\prime}$, there exist for each $\xi\in\mathsf{M}_{q}\Omega$ and $\zeta\in\mathsf{M}_{q}^{\prime}\Omega$ unique operators $x_{\xi}\in\mathsf{M}_{q}$ and $x^{\prime}_{\zeta}\in\mathsf{M}_{q}^{\prime}$ such that $\xi=x_{\xi}\Omega$ and $\zeta=x_{\zeta}^{\prime}\Omega$. Write
@@
Note that the notation $W(\xi)$ is in line with the one introduced above for $\xi\in\mathsf{H}_{\mathbb{R}}$. The operators $W(\xi)$ and $W_{r}(\zeta)$ are respectively called left and right Wick operators. It is known that $\xi_{1}\otimes\ldots\otimes\xi_{n}\in\mathsf{M}_{q}\Omega$ for all $\xi_{i}\in\mathsf{H}_{\mathbb{C}},1\leq i\leq n$ and $\zeta_{1}\otimes\ldots\otimes\zeta_{m}\in\mathsf{M}_{q}^{\prime}\Omega$ for all $\zeta_{j}\in\mathsf{H}_{\mathbb{C}}\cap\operatorname{Dom}(A^{-1/2}),1\leq j\leq m$. By definition, the operators ${W}(\xi)$ and $W_{r}(\zeta)$ commute for all $\xi\in\mathsf{M}_{q}\Omega$ and $\zeta\in\mathsf{M}_{q}^{\prime}\Omega$. In fact, the commutant of $\mathsf{M}_{q}$ in $B(\mathcal{F}_{q}(\mathsf{H}_{U}))$ is given by
@@
For any vectors $\xi_{1},\ldots,\xi_{n}\in\mathsf{H}_{\mathbb{R}}+i\mathsf{H}_{\mathbb{R}}$, the Wick-formula for operators in $\mathsf{M}_{q}$ is given by
@@
For any vectors $\xi_{1},\ldots,\xi_{n}\in\mathsf{H}_{\mathbb{R}}^{\prime}+i\mathsf{H}_{\mathbb{R}}^{\prime}$, the Wick-formula for operators in $\mathsf{M}_{q}^{\prime}$ is given by
@@
The following statements are either well-known or easy to verify.
@@
$c^{*}(f)(\xi\otimes\eta)=(c^{*}(f)\xi)\otimes\eta+q^{m}\xi\otimes c^{*}(f)\eta$
@@
We end this section by recalling the general structure of a one parameter group of orthogonal operators. Let $\{U_{t}\}_{t\in\mathbb{R}}$ be a strongly continuous orthogonal representation of $\mathbb{R}$ on a real Hilbert space $\mathsf{H}_{\mathbb{R}}$ and let $A$ denote the analytic generator of the complexification of $\{U_{t}\}_{t\in\mathbb{R}}$. Let $\mathsf{H}_{\mathbb{R}}^{ap}$ denote the closed real subspace generated by the eigenvectors of $A$ and let $\mathsf{H}_{\mathbb{R}}^{wm}={\mathsf{H}_{\mathbb{R}}^{ap}}^{\perp}$. The subspaces $\mathsf{H}_{\mathbb{R}}^{ap}$ and $\mathsf{H}_{\mathbb{R}}^{wm}$ are reducing under $A$ and every $\{U_{t}\}_{t\in\mathbb{R}}$, and they are respectively called the almost periodic and weakly mixing part of $\mathsf{H}_{\mathbb{R}}$. Similarly, ${U_{t}}_{|_{\mathsf{H}_{\mathbb{R}}^{ap}}}$ and ${U_{t}}_{|_{\mathsf{H}_{\mathbb{R}}^{wm}}}$ are called the almost periodic and weakly mixing part of $\{U_{t}\}_{t\in\mathbb{R}}$ respectively.